40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Message ID: 1416     Entry time: Thu May 12 01:11:07 2011
Author: Zach 
Type: Laser 
Category: GYRO 
Subject: some real progress 

 [Alastair, Koji, Zach]

We had some progress today, which can be summarized in a few parts:

1. Modulation depth

Either I measured this wrong in the previous entry or this magically got better. Koji and I measured it together and we saw sidebands at the level of ~60 mV to the carrier's ~4 V, giving

Γ = 2 sqrt(60 mV / 4 V) = 0.24 rads  This is GOOD.

2. Spurious AM

Continuing with what Alastair started this morning, we tried to reduce the unwanted 19 MHz junk in the PD signal. Alastair pointed out that there was a reasonable amount of noise even with the shutter closed, suggesting some sort of electrical pickup. We tried making some changes to the distribution of the LO signal to the EOM and found that separating it in space from the REFL PD signals returning to the rack made it much better. We also noted that using an SMA cable (instead of BNC with an adapter) for the final stretch over the table to the EOM reduced the pickup further. Koji suggested that I steal a long SMA cable from the 40m so that I could use it for the whole stretch.

3. LF noise

With these improvements the gyro was easily lockable again, and we measured the noise, which we found to be of the same (terrible) level as the other day. We decided to try installing the Cougars into the REFL PDs to add 10 dB of optical gain. This was in part because we had reduced the optical power from >30 mW to 3 mW and figured we could use the gain up front, and in part because we knew that reducing the optical power was one of the only things we had done between now and when our LF noise was lower. I also changed R23 of PDH box #1437 again, this time from 220 to 1k to reduce its gain by ~5 (enough to counteract the added 10 dB of PD gain plus a little more because we were already having to attenuate the PD signal to keep the loop stable).

The result was that the LF f-1/2 noise was shifted down by a little over a factor of 10 (see below). This suggests that the current limiting noise source is probably in the servo electronics and is made less important by shifting the gain to the PDs. I imagine that increasing the optical power while reducing servo gain further will cause it to go down even more (though we shouldn't do this since we want to keep out power at ~3 mW). I am going to do some LISO modeling to see if the expected electronics noise is a plausible culprit given our current configuration.

gyro_noise_wandwo_cougars_5_11_11.pdfimages.jpeg

ELOG V3.1.3-